# Error-Driven Uncertainty Aware Training

### v1

@@ 4.2.1 Image Recognition Models
We start by reporting in Table 2 the results obtained using four models/datasets for image recognition: ResNet50/ImageNet, Wide-ResNet/Cifar100, ResNet18/Cifar10, and ResNet18/SVHN. Across all baselines and metrics, it is evident that EUAT consistently demonstrates superior performance, outperforming others in 16 out of 24 cases. Notably, in the cases where the baselines are more competitive, the differences are marginal (e.g., the error using EUAT increases by 3.4%, 13.8%, and 14.6% compared to the best baseline training a Wide-ResNet/Cifar100, ResNet18/Cifar10, and ResNet18/SVHN, respectively).
@@
EUAT presents the best uA when training a ResNet50 with ImageNet and a Wide-ResNet with Cifar100, and yields minimal reduction of 0.5% and 0.3% in the uA when training a ResNet18 with Cifar10 and SVHN compared with the best baselines (namely, ensemble and CALS, respectively). Additionally, EUAT consistently outperforms existing approaches in terms of uAUC and the correlation between the residuals of the model and predicted uncertainties. EUAT achieves an improvement on the uAUC by up to 15.1%, 8.9%, 64.3%, 28.1%, 8.5%, and 20.4% compared to CE, calibration, DEUP, deep ensemble, CALS, and CE+PE, respectively. Moreover, the correlation between the model’s residuals and predicted uncertainties improves by 20%, 20.7%, 8.6%, and 11.5% when using EUAT for training ResNet50 with ImageNet, Wide-ResNet with Cifar100, and ResNet18 with Cifar10 and SVHN compared to the best-performing baseline in each scenario.
@@
Next, the effectiveness of EUAT in distinguishing correct predictions from misclassifications based on the predicted uncertainty is assessed using the Wasserstein distance of the uncertainty between the sets of correct and wrong predictions. More in detail, on average across all models/datasets, the Wasserstein distance increases by up to 1.7$\times$, 1.3$\times$, 9.7$\times$, 2.3$\times$, 1.4$\times$, and 2.2$\times$ using EUAT compared to CE, calibration, DEUP, ensemble, CALS, and CE+PE, respectively. At last, we evaluate the impact of using EUAT on the ECE and models’ misclassification rate. EUAT results in lower ECE in two cases (ResNet50/ImageNet and ResNet18/SVHN), while a slight increase is observed in the other two (Wide-ResNet/Cifar100 and ResNet18/Cifar10) Importantly, the error rate remains consistent across all benchmarks for all baselines, with a notable improvement observed in the ResNet50/ImageNet benchmark, showcasing a reduction of 13.2% (compared to the best baseline, namely, a deep ensemble). Furthermore, the large gains of EUAT were obtained using larger models and datasets, where the model is less accurate, and thus, by prioritizing the training of incorrect predictions, the EUAT is able to yield significant improvements.
@@
Further, in Figure 1, we plot the distribution of the normalized uncertainty of the correctly and incorrectly predicted sets using the different baselines. By visualizing these distributions, we verify an improvement in the separation of the uncertainty of these two sets using EUAT (which is confirmed by the computation of the Wasserstein distance in Table 2). These results confirm the trustworthiness of the EUAT to improve the model’s uncertainty to separate accurate and incorrect predictions.
@@ 4.2.2 Binary Classification Problem
Next, we proceed by assessing the effectiveness of the EUAT through a binary classification scenario, wherein high-uncertainty predictions can be inverted to the opposite class. We compared the models obtained using the different baselines and tuned the uncertainty threshold as described in Section 3. Additionally, we conducted an extensive evaluation utilizing supplementary metrics such as F1-score, precision, True Positive Rate (TPR), True Negative Rate (TNR), and error rates concerning the inversion or retention of high uncertainty predictions.
Table 3 presents a comprehensive overview of these performance metrics. Remarkably, the EUAT outperforms other baselines across 9 out of 11 evaluated metrics. More in detail, employing EUAT leads to significant enhancements in the F1-score, precision, and TPR, on average, by up to 9.2%, 11.3%, and 11.3%, respectively, with a corresponding 29.9% reduction in error rates through high uncertainty prediction inversion. Furthermore, our method improves the uA, uAUC, and the correlation with model residuals by up to 9.9%, 52.5%, and 113.3%, while the Wasserstein distance of the uncertainty between the correct and wrong predicted sets is enhanced by up to 17.5$\times$ (averaging at 4.1$\times$). Lastly, it is noteworthy that the ECE achieved using the EUAT aligns closely with other baselines, namely CE, DEUP, Ensemble, and CALS, and reduces on average across baselines by 18.1%, while the misclassification rate lowers on average by 6%. These findings underscore the importance of using EUAT to improve the model quality in a binary classification task.
@@ 4.2.3 Out-Of-Distribution Detection Task
Predicted uncertainty can be leveraged to reject difficult examples with high uncertainty, even when these samples present huge distribution shifts from the dataset used to train. Further, the OOD detection task varies significantly when using different methods and datasets [18]. Thus, in this section, we evaluate the effectiveness of EUAT to detect and reject OOD examples based on the predicted uncertainty. For each baseline, we trained a ResNet18 using the Cifar10 dataset and then tested it using a corrupted version with Gaussian noise of this dataset (called Cifar10-C [17]). For a fair comparison, when calibrating the model and training using DEUP, we use an additional validation set containing 10% of the clean inputs of the original test set.
In Table 4, we compared EUAT against the other baselines considering the previous metrics. Notably, in 83% of the metrics considered, EUAT performs better than the baselines. More in detail, EUAT improves the uA by 9.2%, 21.9%, 22.2%, 0.25%, 9.9%, and 10.7% compared to CE, calibration, DEUP, ensembles, CALS, and CE+PE, respectively, while the uAUC increases by 17.7%, 43.9%, 56.7%, 8.4%, 11.2%, and 20.1%. We also verify an enhancement in the correlation between model residuals and the predicted uncertainty, Wasserstein distance of the uncertainty between the correct and wrong predicted sets, and ECE of 3.8$\times$, 16.9$\times$, 2.8$\times$ on average compared to the other baselines. Finally, all the baselines, except CALS, yielded a model with a larger error rate than EUAT.
@@ 4.2.4 Adversarial Training
Finally, we evaluate the EUAT in adversarial training settings. We opted to exclusively train our models with adversarial examples, utilizing the FGSM [13] to generate perturbations, with a predefined perturbation bound $\epsilon$ set to 4/255. Due to resource constraints and the overhead introduced by adversarial training, we did not deploy the ResNet50/ImageNet benchmark in adversarial settings.
We report the results of the different baselines on standard and adversarial settings in Table 5. Despite the challenges posed by adversarial scenarios, overall, we see similar trends compared to the standard training. As expected the adversarial error increases in all baselines and benchmarks considered. In half of the scenarios/metrics assessed, EUAT outperforms the baselines. Quantitatively, employing EUAT yields an average increase in uAUC of 20.2%, 10.9%, and 17.0%, when training a Wide-ResNet on Cifar100, a ResNet18 on Cifar10, and a ResNet18 on SVHN datasets, respectively. Moreover, while the gains in uA are slightly smaller, they still present significant improvements of 15.2%, 1.4%, and 2.1% across the same models/datasets. Additionally, the error rates across baselines exhibit negligible variance, and we verify a larger separation of the uncertainty of the incorrect and correct predictions when performing AT with EUAT, which highlights the robustness of our method across different tasks and domains, reaffirming its efficacy in the challenging context of adversarial attacks.
@@ 5 Conclusion and Future Work
This paper introduces Error-Driven Uncertainty Aware Training, a novel approach designed to refine the estimation of model uncertainty. EUAT is engineered to achieve two primary objectives: first, to heighten uncertainty when models generate inaccurate predictions, and second, to output low uncertainty when predictions are correct. This dual-purpose strategy is achieved through the usage of two loss functions, which adapt based on whether training examples are correctly or incorrectly predicted by the model. By minimizing uncertainty for accurate predictions and maximizing it for mispredictions while retaining error rates, EUAT aims to enhance model reliability.
We evaluate EUAT against six different baselines and using six metrics, and the results consistently demonstrate EUAT’s superior performance across the majority of the considered cases. Even when faced with competitive baselines, EUAT is still able to achieve comparable performance. Furthermore, we extend our evaluation to encompass diverse problems, including binary classification, out-of-distribution detection, and adversarial training settings. Across all evaluated domains, EUAT demonstrates an enhanced ability to differentiate between erroneous and accurate predictions based on uncertainty levels, thereby increasing model trustworthiness.
Given our current focus on evaluating EUAT solely within classification tasks using image recognition models, we envision broadening our scope to encompass other domains such as regression models or language models for machine translation or summarization, which present new and diverse challenges.

### v2

@@ 4.2.1 Image Recognition Models
We start by reporting in Table 2 the results obtained using four models/datasets for image recognition: ResNet50/ImageNet, Wide-ResNet/Cifar100, ResNet18/Cifar10, and ResNet18/SVHN. Across all baselines and metrics, EUAT consistently demonstrates superior performance, outperforming all the other methods in 16 out of 24 cases. Notably, in the cases where alternative approaches are more competitive than EUAT, the performance differences are marginal (e.g., the error using EUAT increases by 3.4%, 13.8%, and 14.6% compared to the best baseline training a Wide-ResNet/Cifar100, ResNet18/Cifar10, and ResNet18/SVHN, respectively).
@@
EUAT presents the best uA when training a ResNet50 with ImageNet and a Wide-ResNet with Cifar100, and yields a minimal reduction of 0.5% and 0.3% in the uA when training a ResNet18 with Cifar10 and SVHN compared with the best baselines (namely, ensemble and CALS, respectively). Additionally, EUAT consistently outperforms existing approaches in terms of uAUC and the correlation between the residuals of the model and predicted uncertainties. EUAT achieves an improvement on the uAUC by up to 15.1%, 8.9%, 64.3%, 28.1%, 8.5%, and 20.4% compared to CE, calibration, DEUP, deep ensemble, CALS, and CE+PE, respectively. Moreover, the correlation between the model’s residuals and predicted uncertainties improves by 20%, 20.7%, 8.6%, and 11.5% when using EUAT for training ResNet50 with ImageNet, Wide-ResNet with Cifar100, and ResNet18 with Cifar10 and SVHN compared to the best-performing baseline in each scenario.
@@
Next, the effectiveness of EUAT in distinguishing correct predictions from misclassifications based on the predicted uncertainty is assessed using the Wasserstein distance of the uncertainty between the sets of correct and wrong predictions. More in detail, on average across all models/datasets, the Wasserstein distance increases by up to 1.7$\times$, 1.3$\times$, 9.7$\times$, 2.3$\times$, 1.4$\times$, and 2.2$\times$ using EUAT compared to CE, calibration, DEUP, ensemble, CALS, and CE+PE, respectively. At last, we evaluate the impact of using EUAT on the ECE and models’ misclassification rate. EUAT results in lower ECE in two cases (ResNet50/ImageNet and ResNet18/SVHN), while a slight increase is observed in the other two (Wide-ResNet/Cifar100 and ResNet18/Cifar10). Importantly, the error rate remains consistent across all benchmarks for all baselines, with a notable improvement observed in the ResNet50/ImageNet benchmark, showcasing a reduction of 13.2% (compared to the best baseline, namely, a deep ensemble). Furthermore, the large gains of EUAT were obtained using larger models and datasets, where the model is less accurate. In these settings, there exist a relatively larger set of incorrectly classified inputs. This, in turn, allows EUAT to use relatively larger/richer training sets, which, ultimately, increases the effectiveness of the proposed method.
@@
Further, in Figure 1, we plot the distribution of the normalized uncertainty of the correctly and incorrectly predicted sets using the different baselines. By visualizing these distributions, we verify an improvement in the separation of the uncertainty of these two sets using EUAT (which is confirmed by the computation of the Wasserstein distance in Table 2). Overall, these results demonstrate that EUAT significantly enhances the ability to leverage model’s uncertainty to discriminate between correct and incorrect predictions. This ability is a key requirement to build trustworthy AI systems, i.e., to identify in which scenarios the output of an ML model should not be trusted and should, therefore, undergo requiring further scrutiny (e.g., review by human evaluators).
@@ 4.2.2 Binary Classification Problem
Next, we proceed to assess the effectiveness of the EUAT in a binary classification scenario. In these settings, when the model prediction has high uncertainty, one can take that information as a signal to invert the model’s prediction and, thus, output the opposite class. We compared the models obtained using the different baselines and tuned the uncertainty threshold as described in Section 4.1. Additionally, we conducted an extensive evaluation utilizing supplementary metrics such as F1-score, precision, True Positive Rate (TPR), and True Negative Rate (TNR). Further, we report the error rates both when flipping and not flipping high uncertainty predictions.
Table 3 presents a comprehensive overview of these performance metrics. Remarkably, the EUAT outperforms other baselines across 9 out of 11 evaluated metrics. More in detail, when considering the geometric mean of the relative improvement of EUAT with respect to all the considered baselines, EUAT attains 7.4%, 3.2%, and 9.1% enhancements in the F1-score, precision, and TPR, respectively. Interestingly, when flipping highly uncertain predictions, EUAT is the only solution for which the error rate drops significantly when compared to the case of no flipping. This demonstrates that EUAT achieves superior performance with respect to the considered baselines to produce high/low model uncertainty in the presence of wrong/correct predictions. Further, still considering the flipping of high uncertainty prediction, the geometric mean of the relative error rate reduction of EUAT vs. all the considered baselines is 25.6%.
Furthermore, our method improves the uA, uAUC, and the correlation with the model residuals by up to 9.9%, 52.5%, and 113.3%, while the Wasserstein distance of the uncertainty between the correct and wrong predicted sets is enhanced by up to 17.5$\times$ (with the geometric mean of 2.1$\times$ across all benchmarks). Lastly, it is noteworthy that the ECE achieved using EUAT aligns closely with other baselines, namely CE, DEUP, Ensemble, and CALS, and reduces the geometric mean across baselines by 10.7%, while the misclassification rate lowers by 4,4%. These findings highlight the benefits deriving from the use of EUAT to improve the model’s predictive quality in binary classification tasks.
@@ 4.2.3 Out-Of-Distribution Detection Task
In this section, we assess the ability of EUAT to estimate the model’s uncertainty in OOD settings, namely when there are strong distribution shifts with respect to the dataset used to train. These are settings which, in prior works, e.g., [18], have been shown to be very challenging for existing uncertainty estimation methods.
More in detail, in this section, we evaluate the effectiveness of EUAT to detect OOD examples based on the predicted uncertainty. For each baseline, we trained a ResNet18 using the Cifar10 dataset and then tested it using a corrupted version with Gaussian noise of the same dataset (called Cifar10-C [17]). For a fair comparison, all the compared techniques use a validation set of the same size (10%) based on the clean inputs of the original test set.
In Table 4, we compared EUAT against the other baselines considering the aforementioned metrics. Notably, in five out of the six considered metrics, EUAT performs better than the baselines. More in detail, EUAT improves the uA by 9.2%, 21.9%, 22.2%, 0.25%, 9.9%, and 10.7% compared to CE, calibration, DEUP, ensembles, CALS, and CE+PE, respectively, while the uAUC increases by 17.7%, 43.9%, 56.7%, 8.4%, 11.2%, and 20.1%. We also verify an enhancement in the correlation between model residuals and the predicted uncertainty, in the Wasserstein distance of the uncertainty between the correct and wrong predicted sets, and in the ECE of 2.6$\times$, 4.4$\times$, 2.42$\times$ (geometric mean) compared to the other baselines. Finally, all the baselines, except CALS, yielded a model with a larger error rate than EUAT.
@@ 4.2.4 Adversarial Training
Finally, we evaluate EUAT in adversarial training settings. We opted to exclusively train our models with adversarial examples, utilizing FGSM [13] to generate perturbations, with a predefined perturbation bound $\epsilon$ set to 4/255. Due to resource constraints and the overhead introduced by adversarial training, we did not deploy the ResNet50/ImageNet benchmark in adversarial settings.
In Table 5, we report the results of the different baselines when using as test set either clean or adversarial data. Overall, even with adversarial data, we observe similar trends compared to the ones already seen in the standard training (Table 2). As expected the adversarial error increases in all baselines and benchmarks considered. In half of the scenarios/metrics assessed, EUAT outperforms the baselines. Quantitatively, employing EUAT yields an increase in the geometric mean of the uAUC (across all baselines) of 16.3%, 8.3%, and 10.1%, when training a Wide-ResNet on Cifar100, a ResNet18 on Cifar10, and a ResNet18 on SVHN datasets, respectively. Moreover, while the gains in uA are slightly smaller, they still present significant improvements in the geometric mean (with respect to all baselines) of 15.2%, 1.4%, and 2.1% across the same models/datasets. Additionally, the error rates across baselines exhibit small variance. Further, with EUAT, we verify a larger separation of the uncertainty of the incorrect and correct predictions (Wasserstein distance), which highlights the robustness of our method across different tasks and domains, reaffirming its efficacy also in the challenging context of adversarial attacks.
@@ 5 Conclusion and Future Work
This paper introduces Error-Driven Uncertainty Aware Training, a novel approach designed to refine the estimation of model uncertainty. EUAT is engineered to achieve two primary objectives: first, to heighten uncertainty when models generate inaccurate predictions, and second, to output low uncertainty when predictions are correct. This dual-purpose strategy is achieved through the usage of two loss functions, which vary depending on whether training examples are correctly or incorrectly predicted by the model. By minimizing uncertainty for accurate predictions and maximizing it for mispredictions, while striving to minimize error rates, EUAT aims to enhance model’s trustworthiness.
We evaluate EUAT using six different baselines and considering six metrics. Our evaluation encompasses diverse problems, including image recognition, binary classification, out-of-distribution detection, and adversarial training settings. Across all the evaluated domains, EUAT demonstrates an enhanced ability to differentiate between erroneous and accurate predictions based on uncertainty levels (evaluated via the Wasserstein distance), thereby increasing model trustworthiness. More in detail, the experimental results showcase EUAT’s superior performance across the majority of the considered cases. Further, even in settings where EUAT is not the most competitive solution, it still achieves comparable performance to the best performing baselines.
The EUAT method has been designed for classification tasks, and evaluated using image recognition models. In the future, we intend to broaden the applicability of EUAT to additional domains such as regression problems or natural language processing tasks.
{ack}
This work was supported by the Fundação para a Ciência e a Tecnología (Portuguese Foundation for Science and Technology) through the Carnegie Mellon Portugal Program under grant SFRH/BD/151470/2021 via projects with reference UIDB/50021/2020 and C645008882-00000055.PRR, by the NSA grant H98230-23-C-0274, and by the Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS) program, where we used the Bridges-2 GPU and Ocean resources at the Pittsburgh Supercomputing Center through allocation CIS220073, which is supported by National Science Foundation grants #2138259, #2138286, #2138307, #2137603, and #2138296.